# On Erlang mixture approximations for differential equations with distributed time delays

### v1

@@ Abstract
In this paper, we propose a general approach for approximate simulation and analysis of delay differential equations (DDEs) with distributed time delays based on methods for ordinary differential equations (ODEs). The key innovation is that we 1) approximate the kernel by the probability density function of an Erlang mixture and 2) use the linear chain trick to transform the approximate DDEs to ODEs. Furthermore, we prove that an approximation with infinitely many terms converges for continuous and bounded kernels and for specific choices of the coefficients. We compare the steady states of the original DDEs and their stability criteria to those of the approximate system of ODEs, and we propose an approach based on bisection and least-squares estimation for determining optimal parameter values in the approximation. Finally, we present numerical examples that demonstrate the accuracy and convergence rate obtained with the optimal parameters and the efficacy of the proposed approach for bifurcation analysis and Monte Carlo simulation. The numerical examples involve a modified logistic equation and a point reactor kinetics model of a molten salt nuclear fission reactor.
@@ 1 Introduction
In recent decades, there has been a significant interest in distributed time delays, and they have been used to model a variety of processes, e.g., in pharmacokinetics and pharmacodynamics (PK/PD) [21], neuroscience [1], side-effects of chemotherapy [31], the Mackey-Glass system [35, 56], which can describe respiratory and hematopoietic diseases, population models in biology [9], and pollution in fisheries [6]. Mechanical [2] and economic [17] processes have also been considered, and we refer to the book by Kolmanovskii and Myshkis [28] for more examples. Furthermore, many theoretical results have been developed specifically for DDEs with distributed time delays. Rahman et al. [40] studied the stability of networked systems with distributed delays, Yuan and Belair [54] present general stability and bifurcation results, Bazighifan et al. [4] analyze oscillations in higher-order DDEs with distributed delays, Cassidy [8] demonstrate the equivalence between cyclic ordinary differential equations (ODEs) and a scalar DDE with a distributed delay, and Aleksandrov et al. [3] study the stability of systems with state-dependent kernels.
@@
Despite the many applications, there exists far less theory and fewer numerical methods and software for DDEs with distributed time delays than for ODEs. Even for DDEs with absolute delays, there exist many numerical methods [5, 37] and significant amounts of off-the-shelf software, e.g., for simulation [46]. However, some modeling software does contain functionality for distributed time delays, e.g., NONMEM [53] and Phoenix [21], and customized numerical methods have also been proposed. Typically, they 1) approximate the integral in the convolution using a quadrature rule, e.g., a trapezoidal [22] or Gaussian [50] rule, and 2) discretize the differential equations using a one-step method, e.g., a Runge-Kutta method [56, 16], or a linear multistep method [51]. Methods based on splines have also been proposed [57]. However, compared to similar methods for ODEs, approximating the convolution is requiring in terms of computations and memory, and it is not straightforward to choose the time step size adaptively.
@@
Consequently, it is desirable to identify approaches that can directly take advantage of existing theory, methods, and software for ODEs, and for a specific class of kernels, it is possible to transform DDEs with distributed delays to ODEs. Specifically, the transformation is called the linear chain trick (LCT) [33, 38, 47], and it is applicable when the kernel is given by the probability density function of an Erlang distribution (referred to as an Erlang kernel). The LCT has been widely used, and Hurtado and Kirosingh [23] and Hurtado and Richards [24, 25] generalized it to transform a wider range of stochastic mean field models to ODEs. Additionally, both Cassidy et al. [10] and Krzyzanski [30] have proposed to simplify the simulation and analysis of specific DDEs by approximating the involved gamma kernels by a hypoexponential kernel and a truncated binomial series, respectively. However, to the best of the author’s knowledge, there are no general-purpose LCTs or approximations for broader classes of kernels.
@@
Therefore, in this work, we present a general-purpose Erlang mixture approximation for continuous and bounded kernels, which allows us to use the LCT to approximate a large class of DDEs with distributed time delays by ODEs. We prove that the approximation converges, and we compare the stability criteria for the approximate ODEs and the original system. Furthermore, we propose a least-squares approach for determining the Erlang mixture approximation coefficients and rate parameter, and we use a numerical example to demonstrate that such an optimal approximation is more accurate than an approximation which uses theoretical coefficients from the proof. Finally, using two numerical examples, we demonstrate that the proposed approach can be used to approximately simulate and analyze the solution of DDEs with distributed time delays.
@@ 2 Preliminaries
The spaces $\mathbb{R}_{\geq 0}=\{x\in\mathbb{R}|x\geq 0\}$ and $\mathbb{N}_{\geq 0}=\{0,1,\ldots\}$ contain the non-negative real numbers and integers, respectively, and $\mathbb{R}_{>0}=\{x\in\mathbb{R}|x>0\}$ is the space of positive real numbers. We denote by $\odot$ the Hadamard (or elementwise) product (also called the Schur product) [19, 20], i.e., if $c=a\odot b$, then $c_{i}=a_{i}b_{i}$. Furthermore, $\det A$ is the determinant of the matrix argument, $A$, and $\operatorname{diag}a$ denotes a diagonal matrix whose diagonal elements are the elements of the vector (or set) argument, $a$, i.e., if $C=\operatorname{diag}a$, then $C_{ii}=a_{i}$ and $C_{ij}=0$ for $i\neq j$. We will make repeated use of the identity
@@
in the proofs in Section 4.1, and we denote by $[\,\cdot\,]^{+}=\max\{0,\cdot\}$ the positive part of the argument.
@@ 3 System
where $r:\mathbb{R}\rightarrow\mathbb{R}^{n_{z}}$ is the delayed variable, and each element of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ is a regular kernel (see Definition 3.1). Furthermore, $h:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{z}}$ is the memory function. We assume that $f$ and $h$ are differentiable in their arguments, and we refer to the paper by Ponosov et al. [38, Thm. 1] for more details on the existence and uniqueness of solutions to the initial value problem (2)–(3). See also the book by Hale and Lunel [18].
@@
(7) $\det\left(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\alpha(s)\,\mathrm{d}sH\right)=0,$
where $I\in\mathbb{R}^{n_{x}\times n_{x}}$ is an identity matrix. The matrices $F\in\mathbb{R}^{n_{x}\times n_{x}}$, $G\in\mathbb{R}^{n_{x}\times n_{z}}$, and $H\in\mathbb{R}^{n_{z}\times n_{x}}$ are the Jacobians of the right-hand side function and the delay function evaluated in the steady state:
@@
4 Mixed Erlang approximation
In this section, we introduce the Erlang kernel and the Erlang mixture kernel, and we prove that the parameters in the latter can be chosen such that it converges to any regular kernel. Figure 1 shows examples of Erlang kernels and Erlang mixture kernels for different rate parameters and coefficients. For simplicity of the presentation, we only consider scalar kernels in this section.
@@ 4 Mixed Erlang approximation
Figure 1: Erlang kernels for $a=1.5$ (left) and $a=1$ (middle) and $m=0,\ldots,5$, and an Erlang mixture kernel (right) for $a=1$, $M=5$, and the following values of the coefficients. Blue: $c_{m}=1/(M+1)$. Red: $c_{0}=0$ and $c_{m}=1/M$ for $m=1,\ldots,M$. Yellow: $c_{0}=c_{M}=1/2$ and $c_{m}=0$ for $m=1,\ldots,M-1$.
4.1 Convergence
Next, we prove that an Erlang mixture kernel of infinite order converges to any regular kernel as the rate parameter, $a$, goes to infinity, for a specific choice of the coefficients $\{c_{m}\}_{m=0}^{\infty}$, which depends on the kernel that is being approximated. First, we define the Erlang mixture delta family, $\delta_{a}$, which is shown in Figure 2 for different values of $t$ and the rate parameter, $a$. Next, we prove a number of lemmas, and we will draw on the analogy of $\delta_{a}$ to the probability density function of a non-negative random variable, e.g., by determining its mean and variance. Finally, we present a general theorem about the approximation of regular kernels and use it to prove the above statement about the Erlang mixture kernel.
@@ 4.1 Convergence
An Erlang mixture delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with rate parameter $a\in\mathbb{R}_{>0}$ is the piecewise constant function
@@
The mean, $\mu:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with a delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, is given by the integral
@@
The variance, $\sigma^{2}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with a delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, is given by the integral
@@
(19) $\delta_{a}(0,s)$ =$=\begin{cases}a,&s\in[0,1/a),\\ 0,&\mathrm{otherwise}.\end{cases}$
Proof 4.7.
By direct substitution, $\ell_{0}(0)=a$ and $\ell_{m}(0)=0$ for $m>0$. Furthermore, by definition, $[s_{0},s_{1})=[0,1/a)$.
@@
The mean, $\mu:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with an Erlang mixture delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with rate parameter $a$ is given by
@@
The variance, $\sigma^{2}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with an Erlang mixture delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with rate parameter $a$ is given by
@@
Let $\delta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be an Erlang mixture delta family with mean $\mu:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ and rate parameter $a\in\mathbb{R}_{>0}$. Then, the inequality
@@
Proof 4.21.
First, we assume that $t>0$ and bound the integral from above and below using Lemma 4.8 and 4.18:
(42) $1\displaystyle 1-\frac{\sigma^{2}(t)}{\bar{\gamma}^{2}}\leq\int_{\mu(t)-\bar{\gamma}}^{\mu(t)+\bar{\gamma}}\delta_{a}(t,s)\,\mathrm{d}s\leq\int_{[\mu(t)-\gamma]^{+}}^{\mu(t)+\gamma}\delta_{a}(t,s)\,\mathrm{d}s\leq\int_{0}^{\infty}\delta_{a}(t,s)\,\mathrm{d}s=$ 1 .
@@
Next, we prove the main results of this section. Theorem 4.22 is an adaptation of the theorem presented in Section 9.3 of the book by Korevaar [29] regarding the convergence of approximations based on delta families.

### v2

@@ Abstract
In this paper, we propose a general approach for approximate simulation and analysis of delay differential equations (DDEs) with distributed time delays based on methods for ordinary differential equations (ODEs). The key innovation is that we 1) approximate the kernel by the probability density function of an Erlang mixture and 2) use the linear chain trick to transform the approximate DDEs to ODEs. Furthermore, we prove that an approximation with infinitely many terms converges for continuous and bounded kernels and for specific choices of the coefficients. We show that the approximate ODEs can be used to assess the stability of the steady states of the original DDEs and that the solution to the ODEs converges if the kernel is also exponentially bounded. Additionally, we propose an approach based on bisection and least-squares estimation for determining optimal parameter values in the approximation. Finally, we present numerical examples that demonstrate the accuracy and convergence rate obtained with the optimal parameters and the efficacy of the proposed approach for bifurcation analysis and Monte Carlo simulation. The numerical examples involve a modified logistic equation, chemotherapy-induced myelosuppression, and a point reactor kinetics model of a molten salt nuclear fission reactor.
@@ 1 Introduction
In recent decades, there has been a significant interest in distributed time delays, and they have been used to model a variety of processes, e.g., in pharmacokinetics and pharmacodynamics (PK/PD) [24], neuroscience [1], side-effects of chemotherapy [34], the Mackey-Glass system [38, 58], which can describe respiratory and hematopoietic diseases, population models in biology [8, 13], and pollution in fisheries [6]. Mechanical [2] and economic [17] processes have also been considered, and we refer to the book by Kolmanovskii and Myshkis [31] for more examples. Furthermore, many theoretical results have been developed specifically for DDEs with distributed time delays. Rahman et al. [43] studied the stability of networked systems with distributed delays, Yuan and Belair [56] present general stability and bifurcation results, Bazighifan et al. [4] analyze oscillations in higher-order DDEs with distributed delays, Cassidy [7] demonstrate the equivalence between cyclic ordinary differential equations (ODEs) and a scalar DDE with a distributed delay, and Aleksandrov et al. [3] study the stability of systems with state-dependent kernels.
@@
Despite the many applications, there exists far less theory and fewer numerical methods and software for DDEs with distributed time delays than for ODEs. Even for DDEs with absolute delays, there exist many numerical methods [5, 40] and significant amounts of off-the-shelf software, e.g., for simulation [49]. However, some modeling software does contain functionality for distributed time delays, e.g., NONMEM [55] and Phoenix [24], and customized numerical methods have also been proposed. Typically, they 1) approximate the integral in the convolution using a quadrature rule, e.g., a trapezoidal [25] or Gaussian [52] rule, and 2) discretize the differential equations using a one-step method, e.g., a Runge-Kutta method [58, 16] or a linear multistep method [53]. Methods based on splines have also been proposed [59]. However, compared to similar methods for ODEs, approximating the convolution is requiring in terms of computations and memory, and it is not straightforward to choose the time step size adaptively.
@@
Consequently, it is desirable to identify approaches that can directly take advantage of existing theory, methods, and software for ODEs, and for a specific class of kernels, it is possible to transform DDEs with distributed delays to ODEs. Specifically, the transformation is called the linear chain trick (LCT) [36, 41, 50], and it is applicable when the kernel is given by the probability density function of an Erlang distribution (referred to as an Erlang kernel). See also [12] on sufficient and necessary conditions under which DDEs can be transformed to ODEs. The LCT has been widely used, and Hurtado and Kirosingh [26] and Hurtado and Richards [27, 28] generalized it to transform a wider range of stochastic mean field models to ODEs. Additionally, both Cassidy et al. [9] and Krzyzanski [33] have proposed to simplify the simulation and analysis of specific DDEs by approximating the involved gamma kernels by a hypoexponential kernel and a truncated binomial series, respectively, and Guglielmi and Hairer [18] consider exponential mixture approximations and demonstrate their approach for DDEs involving both gamma and Pareto kernels. However, to the best of the author’s knowledge, there are no general-purpose LCTs or approximations for broader classes of kernels.
@@
Therefore, in this work, we present a general-purpose Erlang mixture approximation for continuous and bounded kernels, which allows us to use the LCT to approximate a large class of DDEs with distributed time delays by ODEs. In recent work, this approximation was used to develop an algorithm for identifying kernels in DDEs based on partial measurements of the state [45]. We prove that 1) the mixture approximation converges, 2) the approximate ODEs can be used to assess the stability of the steady states of the original DDEs, and 3) the solution to the ODEs converges to that of the DDEs. Furthermore, we propose a least-squares approach for determining the Erlang mixture approximation coefficients and rate parameter. We use numerical examples to demonstrate that such an optimal approximation is more accurate than an approximation which uses theoretical coefficients from the proof, and we compare with the accuracy of one of the exponential mixture approximations described by Guglielmi and Hairer [18] for the model in [33]. Finally, using two numerical examples, we demonstrate that the proposed approach can be used to approximately simulate and analyze the solution of DDEs with distributed time delays.
@@ 2 Preliminaries
The spaces $\mathbb{R}_{\geq 0}=\{x\in\mathbb{R}|x\geq 0\}$ and $\mathbb{N}_{\geq 0}=\{0,1,\ldots\}$ contain the non-negative real numbers and integers, respectively, and $\mathbb{R}_{>0}=\{x\in\mathbb{R}|x>0\}$ and $\mathbb{R}_{\leq 0}=\{x\in\mathbb{R}|x\leq 0\}$ are the spaces of positive and non-negative real numbers. We denote by $\odot$ the Hadamard (or elementwise) product (also called the Schur product) [22, 23], i.e., if $c=a\odot b$, then $c_{i}=a_{i}b_{i}$. Furthermore, $\det A$ is the determinant of the matrix argument, $A$, and $\operatorname{diag}a$ denotes a diagonal matrix whose diagonal elements are the elements of the vector (or set) argument, $a$, i.e., if $C=\operatorname{diag}a$, then $C_{ii}=a_{i}$ and $C_{ij}=0$ for $i\neq j$. We will make repeated use of the identity
@@
in the proofs in Section 4, and we denote by $\cdot\,]^{+}=\max\{0,\cdot\}$ the positive part of the argument.
@@ 3 System
where $r:\mathbb{R}\rightarrow\mathbb{R}^{n_{z}}$ is the delayed variable, and each element of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ is an exponentially bounded regular kernel (see Definition 3.13.2). Furthermore, $h:\mathbb{R}^{n_{x}}\rightarrow\mathbb{R}^{n_{z}}$ is the memory function. We assume that $f$ and $h$ are differentiable in their arguments and that $r$ is bounded, i.e., that $r(t)\leq K_{r}$ for all $t\in(-\infty,t_{f}]$. We refer to the paper by Ponosov et al. [41, Thm. 1] for more details on the existence and uniqueness of solutions to the initial value problem (2)–(3). See also the book by Hale and Lunel [21].
@@
Definition 3.2.
A scalar-valued kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, is exponentially bounded if there exist $L,\rho\in\mathbb{R}_{>0}$ such that
(5) $\alpha(t)$ $\leq Le^{-\rho t}$
for all $t\in\mathbb{R}_{\geq 0}$.
@@
(8) $P(\lambda)=\det\left(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\alpha(s)\,\mathrm{d}sH\right)=0,$
where $P:\mathbb{C}\rightarrow\mathbb{C}$ is the characteristic function and $I\in\mathbb{R}^{n_{x}\times n_{x}}$ is an identity matrix. The matrices $F\in\mathbb{R}^{n_{x}\times n_{x}}$, $G\in\mathbb{R}^{n_{x}\times n_{z}}$, and $H\in\mathbb{R}^{n_{z}\times n_{x}}$ are the Jacobians of the right-hand side function and the delay function evaluated in the steady state:
@@
4 Erlang mixture approximation
In this section, we introduce the Erlang kernel and the Erlang mixture kernel, and we prove that the parameters in the latter can be chosen such that it converges to any regular kernel. Specifically, in Section 4.1, we express an infinite-order Erlang mixture kernel in terms of a so-called delta family and derive expressions for the coefficients based on the kernel, $\alpha$, to be approximated. Next, in Section 4.2, we use the delta family to prove pointwise convergence of the Erlang mixture approximation. In Section 4.3, we present a previous convergence result by Tijms [51] which is used in Section 4.4 to prove several properties of Erlang mixture kernels. These properties are used to prove that the integral of the absolute error converges to zero in Section 4.5 and to relate the stability analyses of the original DDEs in Section 3 and the approximate ODEs described in Section 5. Figure 1 shows examples of Erlang kernels and Erlang mixture kernels for different rate parameters and coefficients. For simplicity of the presentation, we only consider scalar kernels in this section.
@@ 4 Erlang mixture approximation
Figure 1: Erlang kernels for $a=1.5$ (left) and $a=1$ (middle) and $m=0,\ldots,5$, and an Erlang mixture kernel (right) for $a=1$, $M=5$, and the following values of the coefficients. Blue: $c_{m}=1/(M+1)$ for $m=0,\ldots,M$. Red: $c_{0}=0$ and $c_{m}=1/M$ for $m=1,\ldots,M$. Yellow: $c_{0}=c_{M}=1/2$ and $c_{m}=0$ for $m=1,\ldots,M-1$.
4.1 Erlang mixture delta family
In this section, we introduce the Erlang mixture delta family and use it to derive an identity for an infinite-order Erlang mixture kernel, which is used to prove pointwise convergence in Section 4.2. For a given rate parameter, $a$, the identity also provides theoretical expressions for the coefficients $\{c_{m}\}_{m=0}^{\infty}$, which depend on the kernel $\alpha$ that is being approximated. Furthermore, we prove a number of lemmas about the delta family that will also be used in Section 4.2. We will draw on the analogy of $\delta_{a}$ to the probability density function of a non-negative random variable, e.g., by determining its mean and variance. Throughout this section, the time argument $t$ is a fixed parameter.
@@ 4.1 Erlang mixture delta family
For given $t\in\mathbb{R}_{\geq 0}$, an Erlang mixture delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with rate parameter $a\in\mathbb{R}_{>0}$ is the piecewise constant function
@@
Lemma 4.4.
Let $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be a kernel with an integral of one. Then, for all $a\in\mathbb{R}_{>0}$, an infinite-order Erlang mixture kernel $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ with coefficients given by
(17) $c_{m}$ $=\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s,$ $s_{m}$ $=m\Delta s,$ $\Delta s$ $=1/a,$ $m$ $=0,1,\ldots,$
satisfies the identity
(18) $\hat{\alpha}(t)$ $=\int_{0}^{\infty}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s.$
Proof 4.5.
The identity follows directly from the substitution of the definition of the Erlang mixture delta family in (16):
(19) $\int_{0}^{\infty}\delta_{a}(t,s)\alpha(s)\,\mathrm{d}s$ $=\sum_{m=0}^{\infty}\ell_{m}(t)\int_{s_{m}}^{s_{m+1}}\alpha(s)\,\mathrm{d}s.$
When $M$ is infinite and the coefficients, $\{c_{m}\}_{m=0}^{\infty}$, are given by (17), the expression on the right-hand side is in the form (13) in the definition of the Erlang mixture kernel. Furthermore, since the integral of $\alpha$ is one by assumption, the coefficients sum to one for all $a\in\mathbb{R}_{>0}$:
This completes the proof of the identity (18).
@@
For given $t\in\mathbb{R}_{\geq 0}$, the mean, $\mu:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with a delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, is given by the integral
@@
For given $t\in\mathbb{R}_{\geq 0}$, the variance, $\sigma^{2}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with a delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, is given by the integral
@@
(24) $\delta_{a}(0,s)$ $=\begin{cases}a,&s\in[0,1/a),\\ 0,&\mathrm{otherwise}.\end{cases}$
Proof 4.9.
By direct substitution, $\ell_{0}(0)=a$ and $\ell_{m}(0)=0$ for $m>0$. Furthermore, by definition, $s_{0},s_{1})=[0,1/a)$.
@@
For given $t\in\mathbb{R}_{\geq 0}$, the mean, $\mu:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with an Erlang mixture delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with rate parameter $a$ is given by
@@
For given $t\in\mathbb{R}_{\geq 0}$, the variance, $\sigma^{2}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, associated with an Erlang mixture delta family, $\delta_{a}:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, with rate parameter $a$ is given by
@@
Let $\delta:\mathbb{R}_{\geq 0}\times\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$ be an Erlang mixture delta family with mean $\mu:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, variance $\sigma^{2}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, and rate parameter $a\in\mathbb{R}_{>0}$. Then, the inequality
@@
Proof 4.23.
First, we assume that $t>0$ and bound the integral from above and below using Lemma 4.10 and 4.20:
(47) $1-\frac{\sigma^{2}(t)}{\bar{\gamma}^{2}}\leq\int_{\mu(t)-\bar{\gamma}}^{\mu(t)+\bar{\gamma}}\delta_{a}(t,s)\,\mathrm{d}s\leq\int_{[\mu(t)-\gamma]^{+}}^{\mu(t)+\gamma}\delta_{a}(t,s)\,\mathrm{d}s\leq\int_{0}^{\infty}\delta_{a}(t,s)\,\mathrm{d}s=1.$
@@
4.2 Pointwise convergence
In this section, we prove one of the main results (Theorem 4.26), i.e., that the Erlang mixture approximation converges pointwise to any regular kernel, $\alpha$, as the rate parameter goes to infinity. As the delta family $\delta_{a}$ presented in the previous section depends on both $t$ and $s$ (and not only their difference), the proof requires an adaptation (Theorem 4.24) of the theorem presented in Section 9.3 of Chapter 2 in the book by Korevaar [32] on the convergence of approximations based on delta families. We use the pointwise convergence to prove properties of the Erlang mixture approximation in Section 4.4 and to prove convergence of the integral of the absolute error in Section 4.5.